# Interaction-Aware Model Predictive Decision-Making for Socially-Compliant Autonomous Driving in Mixed Urban Traffic Scenarios

### v1

@@ 3.2 Cost Function Formulation
which computed the target acceleration of the AV. The optimization (7) is solved by a Python API of CasADi (Andersson et al., 2019), where the nonlinear programming solver with an interior point optimizer is applied, which can handle both the state (7c), (7d) and input constraints (7e). With these hard constraints, the safety of the IAMPDM can be formally ensured by strictly enforcing operational boundaries.
@@ 3.3 Integrating Intention Modeling
The intention of the pedestrian to cross the street is derived from their non-verbal communication, which includes eye contact, gestures, and body posture. These explicit communication signals are referred to as intention. Pedestrians use these cues similarly when interacting with human-driven vehicles: They give way with hand gestures or signal their right of way by looking at the human driver. These explicit communication signals can be determined by machine-learning-based/data-driven detection systems (see e.g. Yang et al. (2022)) and should be taken into account to extend the MPC for the decision-making. Therefore, the crossing intention $I_{\mathrm{ped}}(t)$ of the pedestrian is introduced as a function of time and pedestrian behavior, which can vary between 00 and $1$, which can be used for predicting the actions of the pedestrian. Note that the aim of this paper is not the development of such intention-detection algorithms. It is assumed for this work that they are given since intention-detection algorithms can be found in the literature (Razali et al., 2021), (Rasouli et al., 2022), (Azarmi et al., 2023), (Zhou et al., 2024). To integrate the pedestrian’s crossing intention into the MPC formulation, the following two extensions are implemented:
@@
However, these elements can sometimes be contradictory, leading to a traffic standstill. Pedestrians often move quickly in the safe zone and slow down in the near zone. They may sometimes wait in the near zone even if they have the right of way or can cross. In such cases, an automated vehicle would stop and wait for the pedestrian [^4] Considering the critical significance of safety, automated vehicles tend to behave conservatively, leading to more frequent stops.. This results in neither the pedestrian nor the vehicle moving, causing a deadlock. As a solution, a discount function of the intention is proposed motivated by game theoretical applications, see e.g. Marín-Solano (2021):
@@
1 Input: $ped$, $veh$, $I_{\mathrm{ped}}(t_{0})$
2 Output: $veh\_acc$
@@
4if $is\_ped\_passed$ or $is\_veh\_passed$ then
@@
7       if $is\_discounting\_intention$ then
@@ 13       Update vehicle state (4)
14       Update $J_{\mathrm{MPC}}$ (6)
15       $veh\_acc\leftarrow$ Solve (7)
16       return $veh\_acc$
17
Algorithm 1 The IAMPDM Algorithm
Note that the IAMPDM can be handle also. Managing dense pedestrian traffic is indeed a challenge in pedestrian crossing scenarios. In scenarios with many pedestrians crossing simultaneously, the IAMPDM framework can be adapted by interacting with the closest person. Another solution, is that the person with the highest estimated crossing intention is taken into account for the interaction. For both cases, the core working principle of the IAMPDM does not change.
@@ 4 Experiment with the Interaction-Aware Model Predictive Decision-Making
    a rule-based decision-making (RBDM) introduced in Varga et al. (2023a).
The independent variable of the study design was the selection of the negotiation algorithm (MPDM, RBDM, NIA). Note: The Ethical Commission of WIVW approved the study based on Code of Ethics at WIVW – Summary for project proposals and articles/papers.
@@ 4.1 Experimental Simulator Platform
The simulator was set up in an $8\times 4$ meter room to simulate an urban street crossing situation for pedestrians. A screen was placed on one wall of the room, projecting the perspective of a pedestrian. This dynamic view corresponded to that of a pedestrian looking perpendicularly along the street in the crossing direction, see Figure 3. To precisely capture the pedestrian’s position, Vive Trackers were used, with the projected view changing according to the position of the test subjects. The main advantage of this configuration lies in avoiding the motion sickness of a virtual reality system, enabling longer usage and accurately representing natural crossing behavior. In contrast, using a VR headset could cause simulator sickness, and the lack of visual representations of body parts in the virtual environment might lead to unnatural movement behavior. Therefore, we decided against a VR experiment based on these considerations and the advice of the ethical commission. of WIVW.
@@
    A simulation computer with SILAB® [^5] SILAB is a software product of WIVW - Würzburg Institute for Traffic Sciences GmbH (www.wivw.de). software framework,
@@
    A short-distance projector to display the scenario with the automated vehicle,
@@
    A body sensor to track the position and velocity of the test subjects and
@@
    An input joystick for detecting the crossing intention of the test subjects.
The advantage of using a joystick as input for the crossing intention helps to overcome the errors of computer vision algorithms that would be necessary to recognize the crossing intention of the pedestrian. Since this work does not attempt to validate such detection algorithms, the use of the joystick is a reasonable solution. Therefore, the experiment is even more rigorous in its objective of comparing different decision-making algorithms in a closed-loop manner.
Figure 3: The simulator setup. Note that the light was turned off during the experiments, and the room was completely dark.
@@ 4.2 The Setups of the Decision-Making Algorithms
An automated tuning framework for pedestrian-AV interaction is presented in Varga et al. (2023a), which utilizes the optimization
@@
$J_{\mathrm{glob}}=\int_{t_{0}}^{T_{\mathrm{end}}}k_{1}\cdot t+k_{2}\cdot\left|a^{2}_{\mathrm{max,veh}}(t)\right|-k_{3}\cdot\left|d_{\mathrm{min}}\right|+k_{4}\cdot\frac{1}{TTC(t)}\,\mathrm{d}t,$
@@
2while Overaall result not satisfactory do
@@ 5       Testing the decision-making
The RBDM is also tuned with Algorithm 2; the only difference is that the RBDM has fewer parameters.
To mitigate the risk of injury in urban areas, the non-interactive algorithm assumes that pedestrians near the street edge may cross unpredictably at any time. As a result, the vehicle halts and waits for a duration of $t_{\mathrm{NIA}}$, disregarding pedestrian crossing intentions since it assumes the absence of high-level communication detection. If the pedestrian does not cross, the vehicle starts moving again cautiously after $t_{\mathrm{NIA}}$, in accordance with the IAMPDM discount function, which leads to similar wait times.
@@ 4.3 Experiment Design
Since existing research does not address the human-in-the-decision-loop nature of pedestrian-AV interactions in unsignalised crossings, a key challenge was to instruct test subjects in a way that 1) facilitates interaction but 2) does not result in a fixed sequence of human actions that entirely predetermines their behavior. This balance is challenging, as too much freedom leads to high variability, whereas fully predefined behavior compromises the evaluation of the IAMPDM. Forcing the interaction may result in unrealistic scenarios and misleading results.
This challenge was addressed by instructing the test subjects to mimic altering their decisions: They started to cross the street, but they waited on the roadside. Therefore, we defined four scenarios, which describe approximately how the test subjects should behave, but no fixed sequence of human actions is defined, no, crossing time and distance specifications for specific actions are given in advance. The detailed instructions are given in the Appendix. These scenarios are
@@ 4.4 Objective Goals and Evaluation Criteria
    The use of intention-aware decision-making in autonomous vehicles significantly does not alter the crossing times and the criticality of intersection scenarios.
@@
To measure our hypotheses, we used three evaluation metrics defined in Westhofen et al. (2023):
@@
    where $\kappa=0.05$ in order to ensure the numerical stability, in cases of the pedestrian or the vehicle ar at a standstill. From that the average $TTC_{\mathrm{avg}}$ is calculated by
@@
Note hypothesis H1 is evaluated by the objective metrics $TTC$, $DST$ and $T_{\mathrm{end}}$. For the assessment of hypothesis H2 the final questions are used.
5 Results and Discussion
This section presents the results of our experiment and provides an in-depth discussion. Please note that the Delayed Crossing and Delayed Remaining scenarios cause different overall behaviors in the pedestrian and the automated vehicle, making it necessary to assess them separately. For both objective and subjective results, the outliers are removed using the inter-quartile range method, see e.g. Dash et al. (2023).
@@ 5.1 Objective Results
To assess whether these differences between the decision-making algorithms are statistically significant, statistical tests are performed. First, the Kruskal-Wallis test is utilized for evaluating H1, because all three samples are compared, see e.g. Dalgaard (2008). The degrees of freedom of this test are $df=2$, and the significance level is chosen to $\alpha=0.01.$ Its null hypothesis is that there is no difference between the three decision-making algorithms. This hypothesis is declined if $\mathcal{H}\geq\mathcal{X}_{df,\alpha}^{2}$ holds, where $\mathcal{X}^{2}_{df=2,\alpha=0.01}=9.21$.
@@
Since $\mathcal{H}_{i}\geq\mathcal{X}^{2}_{df,\alpha}$ is satisfied for $i=\{Rem_{T_{\mathrm{end}}},Rem_{DST},Rem_{TTC}\}$, we can conclude that the choice of decision-making algorithms has statistically significant effects on $T_{\mathrm{end}}$, $DST$ and $TTC$. In addition, Mann–Whitney–Wilcoxon tests are carried out to compare IAMPDM and RBDM. As it can be seen from Table 5, aside from $TCC$ in Delayed Remaining case, there is no statistical difference between IAMPDM and RBDM.
@@
(c) Average TTCs for Delayed Crossing
@@ 5.2 Subjective Results
For the testing of H2, the results of the final questions are analyzed, for which the test subject Nr. 21 is excluded since the person could not recall which decision-making was which. First, the preferences of the test subject are given in Table 6. It can be seen that most of the test subjects have chosen intention-aware decision-making. The results of Q2 from the final questions are given in Table 7. The raw data of Q2 is included in B. It can be seen that in the subjective assessment, both RBDM and the IAMPDM possess larger mean values than NIA. The Kruskal-Wallis test is conducted to test the difference for statistical significance. The result is
$\mathcal{H}_{\mathrm{Subj}}=14.56,$
which indicates that the choice between intention-aware and non-interactive decision-making algorithms leads to statistically significant differences in the preference of the human test subjects. For a comparison between the IAMPDM and RBDM, a Mann–Whitney–Wilcoxon test is carried out, which yields
$p_{\mathrm{Subj}}=0.154,$
indicating that the difference between IAMPDM and RBDM is not significant.
@@ 5.3 Discussion and Limitations
One of the limitations of our experimental setup lies in the assumption of perfect crossing intention detection via joystick input. While this simplified approach enabled controlled testing, real-world intention recognition involves inherent uncertainties due to sensor noise, occlusions, and ambiguous pedestrian behavior.
@@
It is necessary to highlight that higher TTC values are associated with more natural driving behavior of the AV. This preference is reflected in the test subjects choosing the RBDM and IAMPDM over the more conservative NIA algorithm. Additionally, the test subjects provided insights into the general question What do you think of red/green/blue decision-making in general?, which enhanced our understanding of the proposed algorithms. From the study results, we can draw interesting conclusions. A closer examination of the objective results reveals that the intention-aware decision-making algorithms consistently outperform the NIA in all aspects. However, there is no significant difference between IAMPDM and RB, except in the average TTCs concerning Delayed Remaining (see Table 5). Additionally, we observed sporadic instances of extremely high crossing times for IAMPDM in Figure 5(a). These outliers primarily occurred when pedestrians hesitated mid-crossing, triggering repeated velocity adjustments of the AV. This reactive behavior suggests the need for better anticipation of pedestrian motion variability during the prediction.
@@
For the other cases and measures, no significant differences between IAMPDM and RB were observed in this study. These results suggest that a simple algorithm can achieve outcomes comparable to more complex methods in simple urban traffic scenarios. This implies that distinguishing between simple scenarios and complex scenarios is necessary for the development and validation of interaction-aware algorithms for autonomous vehicles. A simple scenario is defined as dyadic interactions with single pedestrians in controlled crossing situations, characterized by predictable trajectories, as in our experiment setup. Complex scenarios involve multiple interacting agents, mixed traffic modes, and ambiguous navigation goals. This distinction is crucial because while rule-based systems suffice for simple cases, complex scenarios demand the adaptive reasoning capabilities of model-based approaches like IAMPDM. The model (1) is also designed for more complex scenarios. It can extend to situations with multiple interacting agents. This flexibility can make it more applicable to a broader range of pedestrian-vehicle interactions, including those involving dynamic or unpredictable behaviors.
@@
Moreover, the subjective results offer interesting insights from the human-in-the-decision-loop experiment. The test subjects, who preferred the NIA, indicated that a more conservative approach could increase safety and be more favored by pedestrians. They had the comment such
@@
Most of the test subjects liked the RBDM more and criticized the long waiting time of the NIA. An interesting result is that the RBDM leads to better subjective results compared to IAMPDM: Most of the test subjects preferred RBDM over the IAMPDM, see Table 6. However, there is no significant difference in the objective results; see Table 7. Furthermore, most of them were not able not distinguish between RBDM and IAMPDM. They left comments like
@@
The reason for this result is the tuning framework, since the human factors component was tuned for both RBDM and IAMPDM in a similar manner, consequently, the subjective results were also similar. Some general comments were that the test subjects were not sure whether the vehicle recognized them or not. To overcome this, implementing a human-machine interface (HMI) for a vehicle to communicate with pedestrians could significantly enhance trust. When a pedestrian sees a clear signal from the vehicle indicating it has detected them, it reduces uncertainty and potential accidents.
The findings suggest that simply increasing the complexity of intention-aware decision-making algorithms does not necessarily lead to better performance or greater pedestrian preference. Additionally, the conditions under which these more complex and reliable model-based decision-making algorithms can achieve a safe scenario need to be validated in further studies.

### v2

@@ 3.2 Cost Function Formulation
which computes the target acceleration of the AV. The optimization (7) is solved by a Python API of CasADi Andersson et al. [2019], where the nonlinear programming solver with an interior point optimizer is applied, which can handle both the state (7c), (7d) and input constraints (7e). With these hard constraints, the safety of the IAMPDM can be formally ensured by strictly enforcing operational boundaries.
@@ 3.3 Integrating Intention Modeling
The intention of the pedestrian to cross the street is derived from their non-verbal communication, which includes eye contact, gestures, and body posture. These explicit communication signals are referred to as intention. Pedestrians use these cues similarly when interacting with human-driven vehicles: They give way with hand gestures or signal their right of way by looking at the human driver. These explicit communication signals can be determined by machine-learning-based/data-driven detection systems (see e.g. Yang et al. [2022]) and should be taken into account to extend the MPC for the decision-making. Therefore, the crossing intention $I_{\mathrm{ped}}(t)$ of the pedestrian is introduced as a function of time and pedestrian behavior, which can vary between 0 and $1$, which can be used for predicting the actions of the pedestrian. Note that the aim of this paper is not the development of such intention-detection algorithms. It is assumed for this work that they are given since intention-detection algorithms can be found in the literature Razali et al. [2021], Rasouli et al. [2022], Azarmi et al. [2023], Zhou et al. [2024]. To integrate the pedestrian’s crossing intention into the MPC formulation, the following two extensions are implemented:
@@
However, these elements can sometimes be contradictory, leading to a traffic standstill. Pedestrians often move quickly in the safe zone and slow down in the near zone. They may sometimes wait in the near zone even if they have the right of way or can cross. In such cases, an automated vehicle would stop and wait for the pedestrian [^1] Considering the critical significance of safety, automated vehicles tend to behave conservatively, leading to more frequent stops.. This results in neither the pedestrian nor the vehicle moving, causing a deadlock. As a solution, a discount function of the intention is proposed motivated by game theoretical applications, see e.g. Marín-Solano [2021]:
@@
1Input: pedped, vehveh, $I_{\mathrm{ped}}(t_{0})$
2 Output: veh_accveh\_acc
@@
4if is_ped_passedis\_ped\_passed or is_veh_passedis\_veh\_passed then
@@
7 if is_discounting_intentionis\_discounting\_intention then
@@ 13   Update vehicle state (4)
14 $veh\_acc\leftarrow$ Optimze (7)
15 return veh_accveh\_acc
16
Algorithm 1 The IAMPDM Algorithm with MPC and Intention Integration
Note that the main benefit of the IAMPDM is that its core structure remains the same across scenarios with few or many pedestrians. The only difference is the number of pedestrian models that need to be included in the MPC formulation. The complete IAMPDM algorithm is summarized in Algorithm 1.
@@ 4 Experiment with the Interaction-Aware Model Predictive Decision-Making
    a rule-based decision-making (RBDM) introduced inVarga et al. [2023a].
The independent variable of the study design was the selection of the negotiation algorithm (MPDM, RBDM, NIA). For the experiment, N = 24 participants, (12 female and 12 male) were recruited from the WIVW Panel. Their mean age was 47.48 years (SD = 16.87). All participants were over 18 years old and had a valid driver’s license. A compensation of 25 Euros was paid for taking part in the experiment. The Ethical Commission of WIVW approved the study based on Code of Ethics at WIVW – Summary for project proposals and articles/papers.
4.1 Description of the Compared Decision-Making Algorithms
4.1.1 Non-Interactive Algorithm (NIA)
The NIA is a conservative, non-interactive baseline that ignores explicit pedestrian crossing intentions; it applies a simple rule to stop the vehicle whenever a pedestrian is detected in the near or crossing zones (see Figure 3) and the time-to-collision (TTC) falls below a predefined threshold. which can lead to unnecessary stops and reduced traffic-flow efficiency compared to intention-aware methods such as IAMPDM and RBDM.
4.1.2 Rule-Based Decision-Making (RBDM)
The RBDM is a decision-making algorithm that considers the pedestrian’s intention to cross the street. The RBDM is based on a set of rules that determine the vehicle’s behavior based on the pedestrian’s position, speed, and intention to cross the street. The RBDM is designed to stop the vehicle if the pedestrian is in the safe zone and the TTC is less than a predefined threshold, but it also considers the pedestrian’s intention to cross the street. If the pedestrian is in the near zone and has a high intention to cross, the RBDM will slow down the vehicle to allow the pedestrian to cross safely. The RBDM is implemented as a comparison algorithm for the IAMPDM. For more details, see Varga et al. [2023a].
@@ 4.2 Experimental Simulator Platform
The simulator was set up in an $8\times 4$ meter room to simulate an urban street crossing situation for pedestrians. A screen was placed on one wall of the room, projecting the perspective of a pedestrian. This dynamic view corresponded to that of a pedestrian looking perpendicularly along the street in the crossing direction, see Figure 4. To precisely capture the pedestrian’s position, Vive Trackers were used, with the projected view changing according to the position of the test subjects. The main advantage of this configuration lies in avoiding the motion sickness of a virtual reality system, enabling longer usage and accurately representing natural crossing behavior. In contrast, using a VR headset could cause simulator sickness, and the lack of visual representations of body parts in the virtual environment might lead to unnatural movement behavior. Therefore, we decided against a VR experiment based on these considerations and the advice of the ethical commission of WIVW.
@@
    A simulation computer running SILAB® [^2] SILAB is a software product of WIVW - Würzburg Institute for Traffic Sciences GmbH (www.wivw.de) software framework that enables real-time traffic simulation with customizable vehicle behaviors and environmental conditions. The software provides accurate physics-based vehicle dynamics and allows integration of different decision-making algorithms for autonomous vehicles.
@@
    A short-distance projector mounted on the ceiling to display the scenario with the automated vehicle. The projector provides a large field of view (approximately 120 degrees horizontal) at a resolution of 1920x1080 pixels, enabling realistic perception of vehicle approach speeds and distances. The projection is calibrated to match real-world scale ratios.
@@
    A body sensor system consisting of HTC Vive Trackers to track the position and velocity of the test subjects in real-time with sub-centimeter accuracy. The tracking system operates at 90 Hz and provides position data of the test subjects that is synchronized with the simulation environment.
@@
    An input joystick for detecting the crossing intention of the test subjects. The joystick provides an input signal ranging from 0 to 1 that represents the strength of the crossing intention. This direct input method avoids the complexity and potential errors of computer vision-based intention recognition systems while still enabling natural interaction dynamics.
Using a joystick to signal crossing intention avoids computer‑vision errors while preserving natural interaction. Since this work does not attempt to validate such detection algorithms, the use of the joystick is a reasonable solution. Therefore, the experiment is even more rigorous in its objective of comparing different decision-making algorithms in a closed-loop manner.
Figure 4: The simulator setup with the marked street, where the pedestrian crosess. Note that the light was turned off during the experiments, and the room was completely dark. The only light source was the projector. The test subject wore the Vive Tracker on the waist to track their position and velocity. The joystick in the test subject’s hand was used to indicate their crossing intention.
@@ 4.3 The Setups of the Decision-Making Algorithms
An automated tuning framework for pedestrian-AV interaction is presented in Varga et al. [2023a], which utilizes the optimization
@@
$J_{\mathrm{glob}}=\int_{t_{0}}^{T_{\mathrm{end}}}k_{1}\cdot t+k_{2}\cdot\left|a^{2}_{\mathrm{max,veh}}(t)\right|$
$-k_{3}\cdot\left|d_{\mathrm{min}}\right|+k_{4}\cdot\frac{1}{TTC_{\mathrm{MPC}}(t)}\,\mathrm{d}t,$ (12)
@@
2while Overall results not sufficient do
@@ 5   Testing the decision-making
The RBDM is also tuned with Algorithm 2; the only difference is that the RBDM has fewer parameters. To mitigate the risk of injury in urban areas, the non-interactive algorithm assumes that pedestrians near the street edge may cross unpredictably at any time. As a result, the vehicle halts and waits for a duration of $t_{\mathrm{NIA}}$, disregarding pedestrian crossing intentions since it assumes the absence of high-level communication detection. If the pedestrian does not cross, the vehicle starts moving again cautiously after $t_{\mathrm{NIA}}$, in accordance with the IAMPDM discount function, which leads to similar wait times.
@@ 4.4 Experiment Design
Since existing research does not address the human-in-the-decision-loop nature of pedestrian-AV interactions in unsignalized crossings, a key challenge was to instruct test subjects in a way that 1) facilitates interaction but 2) does not result in a fixed sequence of human actions that entirely predetermines their behavior. This balance is challenging, as too much freedom leads to high variability, whereas fully predefined behavior compromises the evaluation of the IAMPDM. Forcing the interaction may result in unrealistic scenarios and misleading results.
This challenge was addressed by instructing the test subjects to mimic altering their decisions: They started to cross the street, but they waited on the roadside. Therefore, we defined four scenarios, which describe approximately how the test subjects should behave, but no fixed sequence of human actions is defined, no crossing time and distance specifications for specific actions are given in advance. The detailed instructions are given in the Appendix. These scenarios are
@@ 4.5 Objective Goals and Evaluation Criteria
    The use of intention‑aware decision‑making does not significantly alter the crossing times and the criticality of intersection scenarios.
@@
To measure our hypotheses, we used three evaluation metrics defined in Westhofen et al. [2023]:
@@
    where $\kappa=0.05$ in order to ensure the numerical stability, in cases of the pedestrian or the vehicle are at a standstill. From that the average $TTC_{\mathrm{avg}}$ is calculated by
@@
Note hypothesis H1 is evaluated by the objective metrics $TTC$, DSTDST and $T_{\mathrm{end}}$. For the assessment of hypothesis H2 the final questions are used.
5 Results
This section presents the results of our experiment and provides an in-depth discussion. Please note that the Delayed Crossing and Delayed Remaining scenarios cause different overall behaviors in the pedestrian and the automated vehicle, making it necessary to assess them separately. For both objective and subjective results, the outliers are removed using the inter-quartile range method, see e.g. Dash et al. [2023].
@@ 5.1 Objective Results
To assess whether these differences between the decision-making algorithms are statistically significant, statistical tests are performed. First, the Kruskal-Wallis test is utilized for evaluating H1, because all three samples are compared, see e.g. Dalgaard [2008]. The degrees of freedom of this test are $df=2$, and the significance level is chosen to $\alpha=0.01.$ Its null hypothesis is that there is no difference between the three decision-making algorithms. This hypothesis is declined if $\mathcal{H}\geq\mathcal{X}_{df,\alpha}^{2}$ holds, where $\mathcal{X}^{2}_{df=2,\alpha=0.01}=9.21$.
@@
Since $\mathcal{H}_{i}\geq\mathcal{X}^{2}_{df,\alpha}$ is satisfied for $i=\{Rem_{T_{\mathrm{end}}},Rem_{DST},Rem_{TTC}\}$, we can conclude that the choice of decision-making algorithms has statistically significant effects on $T_{\mathrm{end}}$, $DST$ and $TTC$. In addition, Mann–Whitney–Wilcoxon tests are carried out to compare IAMPDM and RBDM. As it can be seen from Table 5, aside from $TTC$ in Delayed Remaining case, there is no statistical difference between IAMPDM and RBDM.
@@
(c) Average TTCs for Delayed Remaining
@@ 5.2 Subjective Results
For the testing of H2, the results of the final questions are analyzed, for which the test subject Nr. 21 is excluded since the person could not recall which decision-making was which. First, the preferences of the test subject are given in Table 6. It can be seen that most of the test subjects have chosen intention-aware decision-making. The results of Q2 from the final questions are given in Table 7. The raw data of Q2 is included in B. It can be seen that in the subjective assessment, both RBDM and the IAMPDM possess larger mean values than NIA. The Kruskal-Wallis test is conducted to test the difference for statistical significance. The result is $\mathcal{H}_{\mathrm{Subj}}=14.56,$ which indicates that the choice between intention-aware and non-interactive decision-making algorithms leads to statistically significant differences in the preference of the human test subjects. For a comparison between the IAMPDM and RBDM, a Mann–Whitney–Wilcoxon test is carried out, which yields $p_{\mathrm{Subj}}=0.154,$ indicating that the difference between IAMPDM and RBDM is not significant.
@@ 5.3 Discussion and Limitations
After presenting the results, we discuss the implications and limitations of this study. Firstly, it is necessary to highlight that higher TTC values are associated with natural driving behavior of the AV. This preference is reflected in the test subjects choosing the RBDM and IAMPDM over the more conservative NIA algorithm. Additionally, the test subjects provided insights into the general question What do you think of red/green/blue decision-making in general?, which enhanced our understanding of the proposed algorithms.
From the study results, we can draw interesting conclusions. A closer examination of the objective results reveals that the intention-aware decision-making algorithms consistently outperform the NIA in all aspects. However, there is no significant difference between IAMPDM and RBDM, except in the average TTCs concerning Delayed Remaining (see Table 5), where IAMPDM exhibited slightly superior performance. For the other cases and measures, no significant differences between IAMPDM and RBDM were observed in this study.
Additionally, we observed sporadic instances of extremely high crossing times for IAMPDM in Figure 6(a). These outliers primarily occurred when pedestrians hesitated mid-crossing, triggering repeated velocity adjustments of the AV. This reactive behavior suggests the need for better anticipation of pedestrian motion variability during the prediction, see e.g. Tian et al. [2025a].
Road traffic interactions involve reciprocal behaviors where agents adapt based on mutual observations. While Crosato et al. [2023] and Markkula et al. [2020b] use dynamical models that integrate social awareness for decision -making, our work aligns with their frameworks by modeling pedestrian intentions within an MPC structure.
While our work utilizes TTC as a core interaction metric, it is recognized that pedestrians rely on a multitude of cues, see e. g. Tian et al. [2023]. Furthermore, models like the one proposed by Tian et al. [2025a] validate this multi-cue hypothesis. Our IAMPDM can be extended with these additional cues in future work.
@@
The results suggest that a simple algorithm, like the RBDM, can achieve outcomes comparable to more complex methods in simple urban traffic scenarios. This implies that distinguishing between simple scenarios and complex scenarios is necessary for the validation of interaction-aware algorithms for autonomous vehicles. This distinction is crucial because while rule-based systems like RBDM suffice for simple cases, complex scenarios demand the capabilities of model-based approaches like IAMPDM. RBDM is a heuristic algorithm that adjusts velocity based on TTC and intention thresholds (e.g., yielding if pedestrian intention exceeds 0.5), monitoring pedestrian position and velocity. In contrast, IAMPDM employs a predictive model to optimize vehicle trajectories, considering pedestrian intentions and future states. This allows IAMPDM to anticipate pedestrian actions and plan accordingly, which is particularly beneficial in complex scenarios with multiple agents or ambiguous situations.
@@
Moreover, the subjective results offer interesting insights from the human-in-the-decision-loop experiment. The test subjects who preferred the NIA indicated that a more conservative approach could increase safety and be more favored by pedestrians. They had comments such as:
@@
Most of the test subjects preferred RBDM over the IAMPDM, see Table 6. However, there is no significant difference in the objective results; see Table 7. Furthermore, most of them were not able to distinguish between RBDM and IAMPDM, which are reflected in comments such as
@@
The reason for not detecting a significant difference between RBDM and IAMPDM is assumed to be the tuning framework. Both IAMPDM and RBDM were tuned with the same semi-automated, expert-in-the-loop process, which may reduce separability between methods and introduce bias toward smoothness and comfort. The NIA baseline relies on a waiting-time parameter that influences both safety margins and delay.
Some general comments were that the test subjects were not sure whether the vehicle recognized them or not. To overcome this, implementing an eHMI for a vehicle to communicate with pedestrians could significantly enhance trust. When a pedestrian sees a clear signal from the vehicle indicating it has detected them, it reduces uncertainty and potential accidents.
Results indicate that increased algorithmic complexity does not improve performance or user preference in simple, constrained scenarios. Preassumbly, advanced models like IAMPDM require targeted validation in more complex environments with larger interaction spaces to demonstrate their benefits. Future work will include additional metrics and testing in multi-agent, ambiguous settings to further explore these dynamics.
The resulting practical implications for policy and deployment are that policymakers should set naturalness-aligned safety floors (e.g., min-TTC gaps and deceleration bounds), require conservative fallback when intention certainty is low, and mandate dual-objective calibration with fail-soft timeouts to prevent indecision oscillations.